# OPDR: Order-Preserving Dimension Reduction for Semantic Embedding of Multimodal Scientific Data

### v1

@@ Abstract
One of the most costly operations in scientific data management is searching for the $k$ most similar items (or, $k$-nearest neighbors, KNN) from the database after being provided a new item. The state-of-the-art approach to solving this problem is building an index (e.g., DBSCAN, K-means), such that some tuples can be disqualified at an early stage. However, those existing techniques are inappropriate when the scientific data are multimodal as the measures for different modals are incomparable. The recent advances of multimodal machine learning could potentially overcome the multimodality issue by offering an alternative way for building a semantic index, the so-called embedding vectors mapped from the original multimodal data. However, the resulting (embedding) vectors are usually on the order of hundreds or a thousand of dimensions, which are impractically high for time-sensitive scientific applications.
This work proposes to reduce the dimensionality of the output embedding vectors such that the set of top-$k$ nearest neighbors do not change in the lower-dimensional space, namely Order-Preserving Dimension Reduction (OPDR). In order to develop such an OPDR method, our central hypothesis is that by analyzing the intrinsic relationship among key parameters during the dimension-reduction map, a quantitative function may be constructed to reveal the correlation between the target (lower) dimensionality and other variables. To demonstrate the hypothesis, this paper first defines a formal measure function to quantify the KNN similarity for a specific vector, then extends the measure into an aggregate accuracy of the global metric spaces, and finally derives a closed-form function between the target (lower) dimensionality and other variables. We incorporate the closed-function into popular dimension-reduction methods, various distance metrics, and embedding models. Experiments in multiple scientific data sets show that the proposed OPDR method accurately capture the relationship between the target dimensionality and other variables.
@@ Introduction
Many scientific applications involve heterogeneous data resources and cannot directly apply conventional indexing techniques (e.g., DBSCAN, $k$-means) to accelerate query performance. For example, in material sciences (Rangel DaCosta et al. 2021), each material entity comprises multiple pictures along with textual description. Those textural and pictorial (e.g., stored as numerical matrices) data are usually managed by a scientific database where each type of data is managed by a finer-granularity structure, such as tables in relational databases or blobs in NoSQL databases. The conventional wisdom from a data management point of view to efficiently manage those heterogeneous data is to construct a distinct index for each of those structures (e.g., tables, blobs), which does not necessarily reflect the unified, or semantic, measure of the heterogeneous data.
The emergence of multimodal machine learning (e.g., (Li et al. 2022) provides a plausible method to accelerate the queries over heterogeneous scientific data, thanks to the open-source language models that can be directly used for inference; however, the structure of such models cannot be trivially customized and therefore, the potential high dimensionality of the embedding vectors, usually on the order of hundreds or thousands, leads to a problem known as the “curse of dimensionality.” This issue precludes multimodal machine learning from broad deployment due to the high-performance (i.e., time-sensitive) nature of scientific applications. For example, many language models (BERT  (Devlin et al. 2018), ViT (Dosovitskiy et al. 2020), etc.) generate the embedding vectors of 512 or more dimensions; however, scientific applications (Rangel DaCosta et al. 2021; Wu et al. 2020) require to analyze billions of data (i.e., vectors) on the order of milliseconds. The curse of dimensionality exacerbates the computational complexity, not to mention the additional challenges for model accuracy and generalization.
There is therefore an urgent need to develop new methods to fill the gap between the high dimensionality of embedding vectors generated by pretrained language models and the desired lower dimensionality, and this work focuses on one of the most common analytic tasks in scientific applications—$k$-nearest neighbors (KNN) queries. The KNN query is almost ubiquitous in scientific applications due to the nature of scientific discoveries: In many disciplines, especially natural sciences (Lu et al. 2023; Tshitoyan et al. 2019; Gupta et al. 2022), matching a newly found/constructed item to another well-studied entity in the database is one of the most effective methods to study the new item. If the new method could significantly reduce the performance overhead for performing the KNN queries, many scientific applications would benefit from the new method.
@@
In order to reduce the dimension of high-dimensional embedding vectors in a meaningful sense, this paper first proposes a new notion, namely Order-Preserving Measure (OPM), for the set of nearest neighbors that is preserved during a dimension-reduction map between two metric spaces. In essence, the OPM is formalized as a mathematical function that is well defined over the power-set $\sigma$-algebra over the lower-dimensional metric space. This allows us to assign a numerical (real and non-negative) value to an arbitrary subspace of the low-dimensional space and consequently, apply aggregation to quantify the overall accuracy of the dimension-reduction map in terms of preserving the set of $k$-nearest neighbors of each embedding vector.
Built upon the aforementioned OPM, this paper then constructs a global metric to quantify the closeness of two metric spaces $X$ and $Y$ in terms of, again, the preservation of the set of $k$-nearest neighbors from $X$ to $Y$. The objective of constructing such a global metric is to derive a closed-form function among important environmental parameters, such as the dimensionality of metric spaces, the number of data points, and the expected closeness, or accuracy, between the metric spaces in terms of $k$-nearest neighbors.
The conjectured closed-form function is constructed through multiple observations and qualitative analysis. The expected dimension of the lower-dimensional space is positively correlated to both the space’s dimensionality and cardinality (i.e., the number of embedding vectors), but with different weights. Specifically, we assign significantly higher weight (exponentiation) on the closeness metric (also known as accuracy, which will be formally defined) than that on the cardinality.
In addition to theoretical results about the closed-form function, this paper also demonstrates how to incorporate the order-preserving measure and the closed-form function into the workflow of real-world scientific applications. We will detail the implementation of both the OPM and the closed-form function, and demonstrate that how they work together with other subsystems, such as dimension-reduction methods (PCA, MDS), neural network models (CLIP, ViT, Bert), and distance metrics (Euclidean, cosine, Manhattan). We then evaluate the augmented dimension-reduction system with four multimodal scientific datasets of representative characteristics; our results show that the function accurately captures the intrinsic relationship among the key factors of a dimension-reduction map and therefore, contributes new insights to leveraging advanced AI techniques for complex scientific applications.
@@
In summary, this work makes the following scientific and technical contributions.
  • We present a new notion for dimension reduction, namely Order-Preserving Dimension Reduction (OPDR), which leads to a more flexible metric that suffices to capture the invariant of $k$-nearest neighbors.
  • We construct a closed-form function that captures the intrinsic relationship among the target dimensionality and other factors in the context of OPDR.
  • We implement the proposed OPDR method and demonstrate the effectiveness of OPDR by carrying out an extensive evaluation on real-world multimodal scientific data.
Preliminaries and Related Work
Multimodal Scientific Data
Multimodal scientific data has drawn a lot of research interests. For instance, the MELINDA dataset (Wu et al. 2020) focuses on classifying biomedical experiment methods by utilizing multiple data types. Similarly, a study on explainable AI (Jin, Li, and Hamarneh 2022) evaluates algorithms within the context of multimodal medical imaging tasks, emphasizing the importance of decision-making based on raw scientific data.
Researchers have also trained new large language models specifically designed for scientific data. For example, GIT-Mol (Liu et al. 2024) represents a notable effort, developing a multimodal large language model for molecular science that integrates graph, image, and text data. Additionally, MatSciBERT (Gupta et al. 2022) focuses on creating domain-specific language models tailored to materials science. While these models, along with others such as the unsupervised word embeddings for materials science literature (Tshitoyan et al. 2019), demonstrate the potential of specialized models in capturing domain-specific knowledge, they are often limited in scope, typically focusing solely on text-based data.
Multi-modal data is often highly dimensional, making it crucial to find a trade-off between maintain essential information and reducing the data’s complexity. Therefore, some scientists have proposed effective methods to address this challenge. For instance, the study on Multiscale Feature Extraction and Fusion of Image and Text in VQA (Lu et al. 2023) introduces advanced techniques to fuse multimodal data, allowing for a more integrated analysis of different data types. The research on Semi-Supervised Multi-Modal Learning with Balanced Spectral Decomposition (Hu et al. 2020) focuses on optimizing information preservation through spectral decomposition, which enhances the ability to capture correlations within the data. Additionally, Unsupervised word embeddings in materials science literature (Tshitoyan et al. 2019) demonstrates the transformation of textual information into vectors, providing a compact and efficient representation of large-scale text data.
@@
With the introduction of “Attention is All You Need” (Vaswani et al. 2017), the development of large language models has surged, leading to significant advancements in the field. Following this breakthrough, models like BERT (Devlin et al. 2018) and ViT (Dosovitskiy et al. 2020) have focused on text analysis and image recognition respectively. Particularly noteworthy is the CLIP model (Radford et al. 2021), which integrates both text and image data, marking a significant step forward in multimodal data processing. All kinds of studies that improve CLIP model has been propsed. For instance, TiMix (Jiang et al. 2024), address issues with noisy web-harvested text-image pairs by using mixed data samples, while CLIP-Event (Li et al. 2022) and SoftCLIP (Gao et al. 2024) further enhance the alignment between text and images. In (Zhang et al. 2021), authors proposed to convert images into binary codes (hash codes) that preserve image similarity, making it easier to compare images within large-scale datasets.
@@
This section first presents a new notion, i.e., Order-Preserving Measure (OPM), for the set of $k$-nearest neighbors that are preserved during a map between two metric spaces. Then, a closed-form function is constructed to quantify the relationship among the space dimensionality, the number of data points, and the OPM. Finally, a detailed implementation is presented about how to incorporate the new measure and function into scientific applications.
@@
Intuitively speaking, an order-preserving measure (OPM) provides a metric of the number of nearest neighbors that do not change between two spaces. For example, if the $2$-closest points (we assume that the spaces are metric spaces such that pair-wise distances are well defined) of $every$ point in a metric space $(X,\delta_{X})$ are still the same $2$-closest points (of each point) in a metric space $(Y,\delta_{Y})$, then we say the map $f:X\to Y$ is order-preserving of 2, or $OP_{2}$. We will provide a more formal definition of $OP_{z}$, $z\in\mathbb{Z}^{+}\cup\{0\}$ later; before that, we need to point out a common misunderstanding of this notion regarding inclusiveness, as follows.
@@
Mathematically speaking, a measure is a function, say $\mu$, which maps a subset $E_{i}$ of set $X$ in a well-defined $\sigma$-algebra to a value in the extended real line $[0,\infty]$, such that (i) $\mu(\emptyset)$ = 0 and (ii) $\mu\left(\bigcup_{i=1}^{n}E_{i}\right)=\sum_{i+1}^{n}\mu(E_{i})$. Because we assume there exists a well-defined $\sigma$-algebra over (the set of) embedding vectors, say $X$, $E_{i}$’s will be understood as disjoint subsets of $X$ unless otherwise stated. In fact, we will explicitly construct the $\sigma$-algebra on the target space, $Y$. As a side note, a measure function $\mu$ can be also thought of as a homomorphism between the two monoids (i.e., groups without inverses) $(\mathcal{P}(X),\cup)$ and (+$(\mathbb{R}^{+}\cup\{0\},+)$ with kernel $Ker_{\mu}=\{\emptyset\}\subseteq\mathcal{P}(X)$, where $\mathcal{P}(X)$ denotes the power set of $X$. However, our analysis in the following will not require any group-theoretical results.
The construction of the $\sigma$-algebra of embedding vectors is as follows. Let $k$ denote the target number of preserved $k$-nearest neighbors. Let $Y$ denote the target (i.e., low-dimensional) metric space (we omit the metric functions here) of the dimension-reduction map. In practice, the cardinality of $Y$ is finite (i.e., there is an upper limit of number of embedding vectors in scientific applications), which is of course countable. The $\sigma$-algebra $\mathcal{M}_{Y}$ of $Y$ is simply the power set of all the mapped vectors in $Y$, i.e., $\mathcal{M}_{Y}=\mathcal{P}(Y)$, which is obviously a $\sigma$-algebra on $Y$ because $Y$ is countable and $\mathcal{P}(Y)$ is closed for any countable number of set unions.
After constructing its $\sigma$-algebra, we are ready to define the measure on $\mathcal{M}_{Y}$. Let $E^{Y}_{k,i}$ denote the set of the $k$-nearest neighbors of $y_{i}$, where $y_{i}\in Y$. The same notation can be defined for space $X$, which denotes the original metric space where the original vectors of $y_{i}$ reside, i.e., $x_{i}=f^{-1}(y_{i})$, where $f$ denotes the dimension-reduction function. Let $\mu$ be a function defined as follows,
@@
where $F_{j}\in\mathcal{M}_{Y}$. It should be clear that $\mu\in[0,1]\subset[0,+\infty]$ because $\left|E^{Y}_{k,i}\right|\in[0,k]$, where ||$\cdot$|| ⋅ | denotes the cardinality of a set.
We argue that the function defined in Eq. (1) is a measure on $\mathcal{M}_{Y}$. This means that we need to demonstrate two properties of $\mu$ (on $\mathcal{M}_{Y}$): (i) $\mu(\emptyset)=0$ and (ii) $\mu(\bigcup_{i=1}^{\infty}F_{j})=\sum_{i=1}^{\infty}(\mu(F_{j}))$, $\forall F_{j}\subseteq Y$ and $F_{j}$’s are disjoint. Property (i) is trivially satisfied because when $F_{j}=\emptyset$, we have $\mu(F_{j})=\frac{\left|\emptyset\right|}{k}=\frac{0}{k}=0$. To prove property (ii), it suffices to show that if $F_{j}=F_{1}\cup F_{2}$ and $F_{1}\cap F_{2}=\emptyset$, then $\mu(F_{1}\cup F_{2})=\mu(F_{1})+\mu(F_{2})$; this is because $\mathcal{M}_{Y}$ is finite and the above binary relationship can be extended finite times. To simply the notation, let $E=E^{Y}_{k,i}\cap E^{X}_{k,i}$. It follows that $\mu(F_{j})$ = $\frac{\left|(F_{1}\cup F_{2})\cap E\right|}{k}$ = $\frac{\left|(F_{1}\cap E)\cup(F_{2}\cap E)\right|}{k}$. Now, because $F_{1}\cap F_{2}=\emptyset$, we know $(F_{1}\cap E)\cap(F_{2}\cap E)=\emptyset$. This implies that there is no common element between $(F_{1}\cap E)$ and $(F_{2}\cap E)$, which means the sets are separate components and the cardinality of the union set is simply the addition of cardinality of each component: $\left|(F_{1}\cap E)\cup(F_{2}\cap E)\right|=\left|F_{1}\cap E\right|+\left|F_{2}\cap E\right|$. It follows that $\mu(F_{1}\cup F_{2})=\frac{\left|F_{1}\cap E\right|+\left|F_{2}\cap E\right|}{k}=\frac{\left|F_{1}\cap E\right|}{k}+\frac{\left|F_{2}\cap E\right|}{k}=\mu(F_{1})+\mu(F_{2})$, as desired.
The discussion about preserving the set of $k$-nearest neighbors between metric spaces around Eq. (1) is, in our humble opinion, a tip of the iceberg regarding the neighborhood-preserving dimension-reduction maps. The reasoning is that the $k$-nearest neighbors can be thought of the discrete representation of the neighborhood of the referred point, or vector. If we extend the idea to a continuous counterpart, i.e., given a point $p$ and its neighborhood or an open set around it, say $\epsilon(p)$, the question becomes under what conditions the open set of the mapped vectors, namely $f(p)$, corresponds to the open set of $p$ in a dimension-reduction function $f$. In topology and analysis, such a function carrying the open sets in the forward direction is called an open map, which is unfortunately not the same (in fact, the inverse) condition under which the function is continuous. We leave this as an open question to the community.
@@
The previous section defines an additive measure $\mu$ on the $\sigma$-algebra of the projected space $Y$; This section investigates the relationship between $\mu$ and other parameters, such as the space dimensionality and the number of data points, or space cardinality. The idea is that we will construct a function for the above variables that we conjecture are critical during the dimension reduction, largely based on our experiment observations and mathematical intuition. We will later in the evaluation section verify the constructed function as our hypothesis.
@@
where $m=\left|Y\right|=\left|X\right|$, $y_{j}\in Y$, and $u_{i}(\cdot)$ is the measure on $Y$ over $y_{i}$ as defined in Eq. (1). It is not hard to see that the accuracy $A$ defined above falls within the range $[0,1]$, because $\mu_{i}\left(Y\setminus\{y_{i}\}\right)$ is bounded by $k$ (inclusively).
The parameters in our hypothetical function include the accuracy $A_{k}$ between two metric spaces $X$ and $Y$, the number of data points $m=\left|X\right|=\left|Y\right|$, and the dimensionality of the spaces. In practice, the dimension of the domain space $X$ is determined by the neural network model; therefore, our function would assume that $X$’s dimension is a constant and only involve the dimension of $Y$, denoted by $dim(Y)$. Thus, we expect to construct a function $g$ as follows,
@@
The construction of $g$ is as follows. We conjecture the form of $g$ based on the following observations.
@@
Based on the above discussion, we conjecture that the function $g$ takes the following form

### v2

# Order-Preserving Dimension Reduction for Multimodal Semantic Embedding

@@ Abstract
Searching for the $k$-nearest neighbors (KNN) in multimodal data retrieval is computationally expensive, particularly due to the inherent difficulty in comparing similarity measures across different modalities. Recent advances in multimodal machine learning address this issue by mapping data into a shared embedding space; however, the high dimensionality of these embeddings (hundreds to thousands of dimensions) presents a challenge for time-sensitive vision applications. This work proposes Order-Preserving Dimension Reduction (OPDR), aiming to reduce the dimensionality of embeddings while preserving the ranking of KNN in the lower-dimensional space. One notable component of OPDR is a new measure function to quantify KNN quality as a global metric, based on which we derive a closed-form map between target dimensionality and key contextual parameters. We have integrated OPDR with multiple state-of-the-art dimension-reduction techniques, distance functions, and embedding models; experiments on a variety of multimodal datasets demonstrate that OPDR effectively retains recall high accuracy while significantly reducing computational costs.
@@ Introduction
Multimodal retrieval tasks, especially those involving heterogeneous modalities such as text, image (Zhou and Fan 2015), and audio, rely on diverse data sources where conventional indexing techniques such as DBSCAN or $k$-means become ineffective. Many real-world multimodal systems, such as those in material science (Rangel DaCosta et al. 2021), often consist of multiple images and associated textual descriptions, stored in databases that manage structured (tables) and unstructured (blobs) data separately. Traditional indexing approaches fail to capture the unified semantic similarity between different modalities, requiring an alternative representation that aligns heterogeneous data within a common space.
Multimodal machine learning (Li et al. 2022) provides a means to embed heterogeneous data into a shared vector space (Jia et al. 2021), enabling unified similarity search. However, embeddings from different modalities, such as text and images, are often concatenated, leading to an even higher-dimensional joint representation, which further exacerbates the curse of dimensionality (Aremu, Hyland-Wood, and McAree 2020). Given that embeddings generated by models such as BERT (Devlin et al. 2018) and ViT (Dosovitskiy et al. 2020) already contain hundreds to thousands of dimensions, concatenation leads to prohibitively high-dimensional representations. This dimensionality explosion results in increased storage requirements, computational inefficiency, and degraded nearest-neighbor search performance, particularly in large-scale scientific applications (Wu et al. 2020; Tshitoyan et al. 2019).
Dimensionality reduction offers a potential solution by projecting high-dimensional embeddings into a lower-dimensional space while preserving key relationships. Techniques such as PCA (Hotelling 1933) and ISOMAP (Tenenbaum, de Silva, and Langford 2000) aim to minimize the distortion in inter-point distances, facilitating efficient retrieval in reduced spaces. However, a fundamental challenge in applying dimensionality reduction to multimodal retrieval is the lack of a principled method to determine the optimal target dimensionality. Without an a prior understanding of how well nearest-neighbor structures are preserved post-reduction, practitioners must rely on heuristic or empirical tuning, limiting the robustness of these methods in real-world applications (Zhao et al. 2023). This uncertainty is particularly problematic in tasks such as cross-modal retrieval (Liang et al. 2024), Retrieval-Augmented Generation (RAG) (Lewis et al. 2020), and scientific knowledge discovery (Gupta et al. 2022; Lu et al. 2023), where the integrity of neighborhood relationships directly impacts retrieval effectiveness.
These challenges underscore the necessity of understanding the trade-off between dimensionality reduction and nearest-neighbor preservation. A framework that enables quantitative estimation of neighbor preservation after dimensionality reduction would allow for more effective multimodal retrieval by mitigating the curse of dimensionality while maintaining computational efficiency.
@@
To effectively reduce the dimensionality of high-dimensional embedding vectors while preserving the neighborhood structure, we introduce a mathematically rigorous measure function that quantifies the preservation of nearest neighbors across different metric spaces. This measure, formally defined over the power-set $\sigma$-algebra of the lower-dimensional space, quantifies the degree to which $k$-nearest neighbors are preserved. By enabling a well-defined aggregation over the entire space, this measure provides a concrete way to evaluate the fidelity of dimension-reduction transformations.
Building on this measure, we analyze the quantitative relationship between neighborhood preservation and target dimensionality. We construct a closed-form function that captures the intrinsic dependencies among key parameters, including the original and reduced dimensionality, the number of data points, and the expected neighborhood preservation accuracy. Our analysis reveals that the required lower-dimensional space is positively correlated with both the original dimensionality and dataset cardinality, but their influence varies. In particular, neighborhood preservation accuracy exerts an exponentially stronger effect than dataset size, indicating its dominant role in determining the optimal reduced dimensionality.
In addition to theoretical results, we demonstrate how the proposed measure and closed-form function integrate into practical vision applications. We detail their incorporation into dimensionality reduction pipelines using PCA (Hotelling 1933) and MDS (Kruskal and Wish 1978), evaluate their compatibility with state-of-the-art embedding models such as CLIP (Radford et al. 2021), ViT (Dosovitskiy et al. 2020), and BERT (Devlin et al. 2018), and examine their performance under different distance metrics, including Euclidean, cosine, and Manhattan. Extensive experiments on multimodal scientific datasets and image-text benchmarks validate the accuracy of the closed-form function in capturing the relationship between dimensionality reduction and neighborhood preservation. The results highlight the effectiveness of this framework in mitigating the curse of dimensionality while ensuring robust retrieval performance in vision-based applications.
@@
This work makes the following key contributions:
  • We introduce a mathematically rigorous measure function to quantify the preservation of $k$-nearest neighbors across different dimensional spaces, providing a formal foundation for evaluating dimensionality reduction methods.
  • We derive a closed-form function that quantitatively describes the relationship between neighborhood preservation and target dimensionality, offering a principled approach to determining the optimal reduced dimensionality.
  • We empirically validate the proposed framework across multiple datasets and dimensionality reduction techniques, demonstrating that the closed-form function accurately predicts neighborhood preservation and effectively mitigates the curse of dimensionality.
Background and Related Work
Multimodal Data Analytics
Multimodal data analytics has attracted growing research interest across scientific domains. For example, the MELINDA dataset (Wu et al. 2020) focuses on classifying biomedical experiment methods by utilizing multiple data types. Similarly, a study on explainable AI (Jin, Li, and Hamarneh 2022) evaluates algorithms within the context of multimodal medical imaging tasks, emphasizing the importance of decision-making based on raw scientific data. In parallel, researchers have developed new large language models specifically designed for scientific data. For example, GIT-Mol (Liu et al. 2024) represents a notable effort, developing a multimodal large language model for molecular science that integrates graph, image, and text data. Additionally, MatSciBERT (Gupta et al. 2022) focuses on creating domain-specific language models tailored to materials science. While these models, along with others such as the unsupervised word embeddings for materials science literature (Tshitoyan et al. 2019), demonstrate strong potential in capturing domain-specific knowledge, they are often limited in scope, typically focusing solely on text-based data.
Multimodal data is often highly dimensional, making it crucial to find a trade-off between maintaining essential information and reducing the data’s complexity. Given the high dimensionality of multimodal data, researchers have proposed methods to balance information preservation and complexity reduction. For instance, the study on Multiscale Feature Extraction and Fusion of Image and Text in VQA (Lu et al. 2023) introduces advanced techniques to fuse multimodal data, allowing for a more integrated analysis of different data types. The research on Semi-Supervised Multimodal Learning with Balanced Spectral Decomposition (Hu et al. 2020) focuses on optimizing information preservation through spectral decomposition, which enhances the ability to capture correlations within the data. Additionally, Unsupervised word embeddings in materials science (Tshitoyan et al. 2019) demonstrates the transformation of textual information into vectors, providing a compact and efficient representation of large-scale text data.
@@
With the introduction of “Attention is All You Need” (Vaswani et al. 2017), the development of large language models has surged, leading to significant advancements in the field. Following this breakthrough, models like BERT (Devlin et al. 2018) and ViT (Dosovitskiy et al. 2020) have focused on text analysis and image recognition respectively. Particularly noteworthy is the CLIP model (Radford et al. 2021), which integrates both text and image data, marking a significant step forward in multimodal data processing. Various improvements to the CLIP model have been proposed. For instance, TiMix (Jiang et al. 2024), address issues with noisy web-harvested text-image pairs by using mixed data samples, while CLIP-Event (Li et al. 2022) and SoftCLIP (Gao et al. 2024) further enhance the alignment between text and images.
In (Zhang et al. 2021), authors proposed to convert images into binary codes (hash codes) that preserve image similarity, making it easier to compare images within large-scale datasets.
@@
Multimodal Embedding and Retrieval across Heterogeneous Data
Recent progress in multimodal representation learning has significantly advanced cross-modal retrieval tasks, which aim to identify semantically aligned items across heterogeneous modalities such as text, images, and audio. Traditional vector retrieval methods, including FAISS(Douze et al. 2024), ScaNN(Guo et al. 2020), and HNSW (Malkov and Yashunin 2018), enable efficient similarity search for high-dimensional embeddings. However, as embedding dimensionality increases, these methods face scalability challenges, leading to degraded retrieval efficiency and increased computational cost. While existing approaches focus on scaling similarity search or enhancing retrieval precision, few address the structural integrity of nearest-neighbor relations after dimensionality reduction. To address this limitation, our work introduces Order-Preserving Dimension Reduction (OPDR), which reduces embedding dimensionality while preserving $k$-nearest neighbor (kNN) structure, thereby improving retrieval efficiency in large-scale multimodal datasets.
@@
This section first presents a new notion, i.e., Order-Preserving Measure (OPM), for the set of $k$-nearest neighbors that are preserved during a map between two metric spaces. Then, a closed-form function is constructed to quantify the relationship among the space dimensionality, the number of data points, and the OPM. Finally, a detailed implementation is presented about how to incorporate the new measure and function real-world multimodal retrieval systems.
@@
Informally, an order-preserving measure (OPM) provides a metric of the number of nearest neighbors that do not change between two spaces. For example, if the $2$-closest points (we assume that the spaces are metric spaces such that pair-wise distances are well defined) of everyevery point in a metric space $(X,\delta_{X})$ are still the same $2$-closest points (of each point) in a metric space $(Y,\delta_{Y})$, then we say the map $f:X\to Y$ is order-preserving of 2, or $OP_{2}$. We will provide a more formal definition of $OP_{z}$, $z\in\mathbb{Z}^{+}\cup\{0\}$ later; before that, we need to point out a common misunderstanding of this notion regarding inclusiveness, as follows.
@@
Formally, a measure is a function, say $\mu$, which maps a subset $E_{i}$ of set $X$ in a well-defined $\sigma$-algebra to a value in the extended real line $0,\infty]$, such that (i) $\mu(\emptyset)$ = 0 and (ii) $\mu\left(\bigcup_{i=1}^{n}E_{i}\right)=\sum_{i+1}^{n}\mu(E_{i})$. Because we assume there exists a well-defined $\sigma$-algebra over (the set of) embedding vectors, say $X$, $E_{i}$’s will be understood as disjoint subsets of $X$ unless otherwise stated. In fact, we will explicitly construct the $\sigma$-algebra on the target space, $Y$. As a side note, a measure function $\mu$ can be also thought of as a homomorphism between the two monoids (i.e., groups without inverses) $(\mathcal{P}(X),\cup)$ and $(\mathbb{R}^{+}\cup\{0\},+)$ with kernel $Ker_{\mu}=\{\emptyset\}\subseteq\mathcal{P}(X)$, where $\mathcal{P}(X)$ denotes the power set of $X$. However, our analysis in the following will not require any group-theoretical results.
The construction of the $\sigma$-algebra of embedding vectors is as follows. Let $k$ denote the target number of preserved $k$-nearest neighbors. Let $Y$ denote the target (i.e., low-dimensional) metric space (we omit the metric functions here) of the dimension-reduction map. In practice, the cardinality of $Y$ is finite (i.e., there is an upper limit of number of embedding vectors in vision applications), which is of course countable. The $\sigma$-algebra $\mathcal{M}_{Y}$ of $Y$ is simply the power set of all the mapped vectors in $Y$, i.e., $\mathcal{M}_{Y}=\mathcal{P}(Y)$, which is obviously a $\sigma$-algebra on $Y$ because $Y$ is countable and $\mathcal{P}(Y)$ is closed for any countable number of set unions.
After constructing its $\sigma$-algebra, we are ready to define the measure on $\mathcal{M}_{Y}$. Let
$E^{Y}_{k,i}$ denote the set of the $k$-nearest neighbors of $y_{i}$, where $y_{i}\in Y$. The same notation can be defined for space $X$, which denotes the original metric space where the original vectors of $y_{i}$ reside,
i.e., $x_{i}=f^{-1}(y_{i})$, where $f$ denotes the dimension-reduction function. Let $\mu$ be a function defined as follows,
@@
where $F_{j}\in\mathcal{M}_{Y}$. It should be clear that $\mu\in[0,1]\subset[0,+\infty]$ because $\left|E^{Y}_{k,i}\right|\in[0,k]$, where $|\cdot|$ denotes the cardinality of a set.
We argue that the function defined in Eq. (1) is a measure on $\mathcal{M}_{Y}$. This means that we need to demonstrate two properties of $\mu$ (on $\mathcal{M}_{Y}$): (i) $\mu(\emptyset)=0$ and (ii) $\mu(\bigcup_{i=1}^{\infty}F_{j})=\sum_{i=1}^{\infty}(\mu(F_{j}))$, $\forall F_{j}\subseteq Y$ and $F_{j}$’s are disjoint. Property (i) is trivially satisfied because when $F_{j}=\emptyset$, we have $\mu(F_{j})=\frac{\left|\emptyset\right|}{k}=\frac{0}{k}=0$. To prove property (ii), it suffices to show that if $F_{j}=F_{1}\cup F_{2}$ and $F_{1}\cap F_{2}=\emptyset$, then $\mu(F_{1}\cup F_{2})=\mu(F_{1})+\mu(F_{2})$; this is because $\mathcal{M}_{Y}$ is finite and the above binary relationship can be extended finite times. To simplify the notation, let $E=E^{Y}_{k,i}\cap E^{X}_{k,i}$. It follows that $\mu(F_{j})$ = $\frac{\left|(F_{1}\cup F_{2})\cap E\right|}{k}$ = $\frac{\left|(F_{1}\cap E)\cup(F_{2}\cap E)\right|}{k}$. Now, because $F_{1}\cap F_{2}=\emptyset$, we know $(F_{1}\cap E)\cap(F_{2}\cap E)=\emptyset$. This implies that there is no common element between $(F_{1}\cap E)$ and $(F_{2}\cap E)$, which means the sets are separate components and the cardinality of the union set is simply the addition of cardinality of each component: $\left|(F_{1}\cap E)\cup(F_{2}\cap E)\right|=\left|F_{1}\cap E\right|+\left|F_{2}\cap E\right|$. It follows that $\mu(F_{1}\cup F_{2})=\frac{\left|F_{1}\cap E\right|+\left|F_{2}\cap E\right|}{k}=\frac{\left|F_{1}\cap E\right|}{k}+\frac{\left|F_{2}\cap E\right|}{k}=\mu(F_{1})+\mu(F_{2})$, as desired.
The discussion about preserving the set of $k$-nearest neighbors between metric spaces around Eq. (1) is, in our humble opinion, a tip of the iceberg regarding the neighborhood-preserving dimension-reduction maps. The reasoning is that the $k$-nearest neighbors can be thought of the discrete representation of the neighborhood of the referred point, or vector. If we extend the idea to a continuous counterpart, i.e., given a point $p$ and its neighborhood or an open set around it, say $\epsilon(p)$, the question becomes under what conditions the open set of the mapped vectors,
namely $f(p)$, corresponds to the open set of $p$ in a dimension-reduction function $f$. In topology and analysis, such a function carrying the open sets in the forward direction is called an open map, which is unfortunately not the same (in fact, the inverse) condition under which the function is continuous. We leave this as an open question to the community.
@@
The previous section defines an additive measure $\mu$ on the $\sigma$-algebra of the projected space $Y$; This section investigates the relationship between $\mu$ and other parameters, such as the space dimensionality and the number of data points, or space cardinality. The idea is to construct a function for the above variables which we hypothesize to be critical for dimensionality reduction, largely based on our empirical observations and mathematical intuition. We will later verify the constructed function as a working hypothesis in the evaluation section.
@@
where $m=\left|Y\right|=\left|X\right|$, $y_{j}\in Y$, and $\mu_{i}(\cdot)$ is the measure on $Y$ over $y_{i}$ as defined in Eq. (1). It is not hard to see that the accuracy $A$ defined above falls within the range $[0,1]$, because $\mu_{i}\left(Y\setminus\{y_{i}\}\right)$ is bounded by $k$ (inclusively).
The parameters in our hypothetical function include the accuracy $A_{k}$ between two metric spaces $X$ and $Y$, the number of data points $m=\left|X\right|=\left|Y\right|$, and the dimensionality of the spaces. In practice, the dimension of the domain space $X$ is determined by the neural network model; therefore, our function would assume that $X$’s dimension is a constant and only involves the dimension of $Y$, denoted by $dim(Y)$. Thus, we expect to construct a function $g$ as follows,
@@
The construction of $g$ is based on the following observations:
@@
Based on the above discussion, we postulate that the function $g$ takes the following form